# Insulator phases of Bose–Fermi mixtures induced by intraspecies next-neighbor interactions

### v2

@@ 3.3 Four fifths fermionic filling $(\rho_{\rm{F}}=4/5)$
We show the corresponding density profiles for the insulator at $\rho_{\rm{B}}=7/15$ in Fig. 8(b). Here we observe the characteristic CDW coupling between spin-up fermions and bosons with seven particles per period of each coupled component, indicating that it also corresponds to a SSCDW phase. The insulator at $\rho_{\rm{B}}=1/3$ also follows an analogous behavior with the corresponding periodicity, not shown here. Moreover, by looking at the bosonic correlation function $\expectationvalue{\hat{b}^{\dagger}_{x}\hat{b}_{x+y}}$  [^1] The correlation functions $\expectationvalue{\hat{b}^{\dagger}_{x}\hat{b}_{x+y}}$ are calculated by averaging over $x$ to remove finite size effects due to the CDW oscillations in the system. in Fig. 9(a) for the imbalanced case from Fig. 8(a) at $\rho_{\rm B}=2/5$ we see a potential decay indicating its SF nature until an edge effect appears as a sudden decrease in the correlation function, while the SSCDW plateaus at $\rho_{\rm B}=1/3$ and $7/15$ exhibit an exponential decay, characteristic of an insulator state; hence the SSCDW states are separated by a superfluid phase [93, 94, 67].
@@
Figure 9: (a) Bosonic correlation function $\expectationvalue{\hat{b}^{\dagger}_{x}\hat{b}_{x+y}}$ averaged over $x$ as a function of $y$ for a lattice of size $L=135$ at the fermionic densities $\rho_{\rm F}^{\uparrow}=7/15$ and $\rho_{\rm F}^{\downarrow}=1/3$ and different bosonic densities, $\rho_{\rm B}=1/3$ (light blue), $2/5$ (orange), and $7/15$ (dark blue). We use a log-log scale to emphasize the superfluid or insulating behavior of the correlation functions and consider the interaction parameters $U_{FF}=4$, $U_{BF}=6$, and $V_{FF}=1$. (b) Spin-up fermionic density $\rho_{\rm F}^{\uparrow}$ vs. the corresponding fermionic chemical potential $\mu_{\rm F}^{\uparrow}$ at the thermodynamic limit with interactions $U_{\rm FF}=4$, $U_{\rm BF}=6$, $V_{\rm FF}=1$ and the rest of the densities given by $\rho_{\rm B}=1/3$ and $\rho_{\rm F}^{\downarrow}=1/3$. The inset shows the extrapolations of the fermionic charge gap $\Delta_{\rm F}^{\uparrow}$ for $\rho_{\rm F}^{\uparrow}=1/3$, $7/15$, and $2/3$. The lines are visual guides, while the points correspond to DMRG results
@@
Now we focus on one of the split SSCDW insulators, specifically $\rho_{\rm B}=1/3$, and show that the uncoupled fermionic spin component is not insulating. For this, we fix $\rho_{\rm B}=\rho_{\rm F}^{\downarrow}=1/3$ and study the dependence of $\rho_{\rm F}^{\uparrow}$ on its corresponding chemical potential $\mu_{\rm F}^{\uparrow}$, which we calculate in an analog way as (5). This graph, shown in Fig. 9(b), exhibits the MMI and SSMI phases at $\rho_{\rm F}^{\uparrow}=1/3$, $2/3$, respectively, with corresponding gaps $\Delta_{\rm{F},\uparrow}^{\rho_{\rm{F}}^{\uparrow}=1/3}=0.93$ and $\Delta_{\rm{F},\uparrow}^{\rho_{\rm{F}}^{\uparrow}=2/3}=2.12$. Nevertheless, for $\rho_{\rm F}^{\uparrow}=7/15$, the density at which the bosonic charge gap is nonzero (see Fig. 8(a)), the spin-up fermionic gap is closed at the thermodynamic limit, emphasized by the inset of Fig. 9(a). This indicates that the spin-up fermions are in a gapless phase inside this SSCDW insulator.
@@
After establishing the spin-selective character of the SSCDW phase, we propose a corresponding order parameter, $\mathcal{O}^{\rm SSCDW}_{\sigma,\rho}$, as an extension of (8)
@@
where we specify the spin as $\sigma=\uparrow$, $\downarrow$  [^2] The SSCDW phase exhibits the same finite-size effect as the MCDW phase where one has to measure its properties with one boson less than the expected, then we use the same procedure from Sec. 3.2 for the calculation of the bosonic gaps, density profiles, and order parameters.. In Fig. 7(b), we show $\mathcal{O}^{\rm SSCDW}_{\sigma,\rho=2/5}$ for a range of $V_{\rm FF}$ values with the parameters of the balanced case from Fig. 7(a), for which we use either $\sigma=\uparrow$ or $\downarrow$ and $\rho_{\rm F}^{\sigma}=\rho_{\rm F}/2$. We note that after the critical value ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=2/5}\approx 0.1$, both the order parameter and the corresponding gap (Fig. 7(a)) increase from zero. For an imbalanced case, we checked that the definition (9) also works as an order parameter using the corresponding coupled fermionic component. Inside the balanced SSCDW phase, half of the fermions are uncoupled from the insulator structure. This decreases its stability with higher $V_{\rm FF}$, as seen in a reduction of both the gap and the order parameter in Fig. 7. For large enough $V_{\rm FF}$, we expect this behavior to break the insulator and transition the ground state to a SF phase. Even then, the SSCDW phase appears only for $\rho_{\rm{B}}+\rho_{\rm{F}}\geq 1$, since in this region the chain gets saturated with all of the carriers, which forces the coupling to only contain one type of fermion and not both, in contrast to what happens with the MCDW. This condition is also independent of the spin-population imbalance, since it only deals with the complete fermionic and bosonic densities.
@@ 3.4 Phase diagram for nonzero spin-population imbalance
According to the previous sections, there are two spin-selective phases in the present non-local mixture model, so it is of great value to analyze how the phase diagram from Fig. 1(a) changes as we increase the spin-population imbalance. This is depicted in Fig. 10, where we exhibit the phase diagram for $I=0$, $1/6$, $1/2$, $1$.
@@
Here, we observe that the phase space configuration of the MMI (dark blue) and MCDW (brown) insulators is invariant under the change of the spin-population imbalance, which further emphasizes its spin-independent nature. On the other hand, as we increase $I$, we can see the splitting of the spin-selective phases. In the case of the SSMI (dark green), the separation grows until the spin-up insulator follows $\rho_{\rm{B}}=1$, while the other one matches the MMI line. For the SSCDW (light green), there is an additional constraint, $\rho_{\rm{B}}+\rho_{\rm{F}}\geq 1$; hence as the imbalance increases, the spin-down phase shortens until it vanishes at $I=1$, while the spin-up phase follows the relation $\rho_{\rm{B}}-\rho_{\rm{F}}=0$. In the case of the ICDW (purple line), the condition $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$ limits its phase space line as $I$ grows until it converges at the point $\rho_{\rm{B}}=\rho_{\rm{F}}=1/2$. Hence for polarized fermions at half-filling the three different insulators found in this study follow the single condition $\rho_{\rm{B}}-\rho_{\rm{F}}=0$. This emphasizes how a study with spinor fermions can show different emergent behaviors that can be elusive in an investigation with only scalar particles.
@@ 4 Bosonic next-neighbor interactions
Up to this point, we have only considered nonzero next-neighbor fermionic interactions, so a good question is what happens when the non-local interaction is between the bosons only. In order to answer this, in Fig. 11 we plot the $\rho_{\rm{B}}-\mu_{\rm{B}}$ graph for $\rho_{\rm{F}}=2/5$, $2/3$, $4/5$, keeping the local interactions constant at $U_{\rm{FF}}=4$ and $U_{\rm{BF}}=6$ while turning on the next-neighbor bosonic interaction $V_{\rm{BB}}$ for values in the range $0\leq V_{\rm{BB}}\leq 6$.
@@
Figure 11: Bosonic density $\rho_{\rm{B}}$ vs. bosonic chemical potential $\mu_{\rm{B}}$ at the thermodynamic limit for fermionic densities (a) $\rho_{\rm{F}}=2/3$, (b) $\rho_{\rm{F}}=2/5$, and (c) $\rho_{\rm{F}}=4/5$. The local repulsion couplings are $U_{\rm{FF}}=4$, and $U_{\rm{BF}}=6$, and we compare the behavior with and without next-neighbor interaction between bosons using $V_{\rm{BB}}=0$, $2$, $3$, $6$. Each inset shows the charge gap $\Delta_{\rm{B}}$ as a function of the inverse of the system size $L$ for the corresponding nonzero $V_{\rm{BB}}$ at characteristic bosonic densities $\rho_{\rm{B}}$. The points correspond to DMRG results and the lines are visual guides.
@@
For $\rho_{\rm{F}}=2/3$ (Fig. 11(a)) and $V_{\rm{BB}}=2$, we once again find three plateaus corresponding to the MMI at $\rho_{\rm{B}}=1/3$, the SSMI at $\rho_{\rm{B}}=2/3$, and the ICDW at $\rho_{\rm{B}}=1/2$, with charge gaps $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/3}=1.27$, $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/3}=2.51$, and $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=1.18$, respectively. In this case, the main difference concerning $V_{\rm{FF}}\neq 0$ is that the plateaus of the MMI increase slightly, instead of decreasing, as happens in Fig. 4(a), which shows that the charge gap has a dependence on the associated particle statistics of the non-local interactions. Apart from that, for $\frac{1}{2}\leq\rho_{\rm{F}}\leq\frac{2}{3}$ we found that just as in Fig. 11(a), the phase diagram for fermionic next-neighbor interactions is the same for bosonic next-neighbor interactions (See Fig. 1).
@@
On the other hand, at $\rho_{\rm{F}}=2/5$ (Fig. 11(b)) we do observe an additional incompressible phase besides the ones in Fig. 4(a) when we turn on the non-local interaction to $V_{\rm{BB}}=6$. This new plateau is located at $\rho_{\rm{B}}=1/2$ and corresponds to the ICDW with charge gap $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=6.64$. Here, since the bosons possess the non-local interactions, they can form the CDW structure even in the absence of fermions, so in this case the ICDW appears for any fermionic filling, as can be seen in Fig. 1(b). Apart from this insulator, we still find the MMI and SSMI at $\rho_{\rm{B}}=3/5$ and $\rho_{\rm{B}}=4/5$ with gaps $\Delta_{\rm{B}}^{\rho_{\rm{B}}=3/5}=2.09$ and $\Delta_{\rm{B}}^{\rho_{\rm{B}}=4/5}=2.31$, respectively. For this instance, both of them increase their stability with higher $V_{\rm{BB}}$. In Sec. 3.2 for $V_{\rm{FF}}\neq 0$ we found that the non-local interactions drastically reduced the spin-selective phase gap, since the free fermions are enhanced by this interaction and hence can disturb the insulating structure. In this case, only the bosons have non-local interactions, and because of that the stability of the phase is not damped. The MCDW phase still appears in the phase diagram, but with a particularly small charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/5}=0.26$. Considering the high value of the bosonic interaction in this case, this suggests that the MCDW structure is more favorable for $V_{\rm{FF}}\neq 0$.
@@
Finally, in Fig. 11(c) we show the case for $\rho_{\rm{F}}=4/5$, which has results analogous to the previous two fermionic densities. The ICDW appears with a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=3.16$, as expected from Fig. 1(b). For the MMI and SSMI, their charge gaps increase to $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/5}=1.16$ and $\Delta_{\rm{B}}^{\rho_{\rm{B}}=3/5}=2.88$, respectively, in the same way as happened for Fig. 11(a). On the contrary, the SSCDW arises with a small charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/5}=0.26$, similar to the MCDW in Fig. 11(b).

### v3

@@ 3.3 Four fifths fermionic filling $(\rho_{\rm{F}}=4/5)$
We show the corresponding density profiles for the insulator at $\rho_{\rm{B}}=7/15$ in Fig. 8(b), here we observe the characteristic CDW coupling between spin-up fermions and bosons with seven particles per period of each coupled component, indicating that it also corresponds to a SSCDW phase. The insulator at $\rho_{\rm{B}}=1/3$ also follows an analogous behavior with the corresponding periodicity, not shown here. Moreover, by looking at the bosonic correlation function $\expectationvalue{\hat{b}^{\dagger}_{x}\hat{b}_{x+y}}$  [^1] The correlation functions $\expectationvalue{\hat{b}^{\dagger}_{x}\hat{b}_{x+y}}$ are calculated by averaging over $x$ to remove finite size effects due to the CDW oscillations in the system. in Fig. 9(a) for the imbalanced case from Fig. 8(a) at $\rho_{\rm B}=2/5$ we see a potential decay indicating its SF nature until an edge effect appears as a sudden decrease in the correlation function, while the SSCDW plateaus at $\rho_{\rm B}=1/3$, and $7/15$ possess an exponential decay, characteristic of an insulator state, hence the SSCDW states are separated by a superfluid phase.
@@
Figure 9: (a) Bosonic correlation function $\expectationvalue{\hat{b}^{\dagger}_{x}\hat{b}_{x+y}}$ averaged over $x$ as a function of $y$ for a lattice of size $L=135$ at the fermionic densities $\rho_{\rm F}^{\uparrow}=7/15$, $\rho_{\rm F}^{\downarrow}=1/3$ and different bosonic densities $\rho_{\rm B}=1/3$ (light blue), $2/5$ (orange) and $7/15$ (dark blue). We use a log-log scale to emphasize the superfluid or insulating behavior of the correlation functions and consider the interaction parameters $U_{FF}=4$, $U_{BF}=6$ and $V_{FF}=1$. (b) Spin-up fermionic density $\rho_{\rm F}^{\uparrow}$ vs the corresponding fermionic chemical potential $\mu_{\rm F}^{\uparrow}$ at the thermodynamic limit with interactions $U_{\rm FF}=4$, $U_{\rm BF}=6$, $V_{\rm FF}=1$ and the rest of the densities given by $\rho_{\rm B}=1/3$ and $\rho_{\rm F}^{\downarrow}=1/3$. The inset shows the extrapolations of the fermionic charge gap $\Delta_{\rm F}^{\uparrow}$ for $\rho_{\rm F}^{\uparrow}=1/3$, $7/15$ and $2/3$. The lines are visual guides, while the points correspond to DMRG results
@@
Now we focus on one of the splitted SSCDW insulators, specifically $\rho_{\rm B}=1/3$, and show that the uncoupled fermionic spin component is not insulating. For this we fix $\rho_{\rm B}=\rho_{\rm F}^{\downarrow}=1/3$ and study the dependence of $\rho_{\rm F}^{\uparrow}$ on its corresponding chemical potential $\mu_{\rm F}^{\uparrow}$, which we calculate in an analog way as (7). This graph, shown in Fig. 9(b), exhibits the MMI and SSMI phases at $\rho_{\rm F}^{\uparrow}=1/3$, $2/3$, respectively with corresponding gaps $\Delta_{\rm{F},\uparrow}^{\rho_{\rm{F}}^{\uparrow}=1/3}=0.93$ and $\Delta_{\rm{F},\uparrow}^{\rho_{\rm{F}}^{\uparrow}=2/3}=2.12$, nevertheless for $\rho_{\rm F}^{\uparrow}=7/15$, density at which the bosonic charge gap is nonzero (see Fig. 8(a)), the spin-up fermionic gap is closed in the thermodynamic limit, emphasized by the inset of Fig. 9(a). This indicates that the spin-up fermions are in a gapless phase inside this SSCDW insulator.
@@
After establishing the spin-selective character of the SSCDW phase, we propose a corresponding order parameter $\mathcal{O}^{\rm SSCDW}_{\sigma,\rho}$ as an extension of (10)
@@
where we specify the spin as $\sigma=\uparrow$, $\downarrow$  [^2] The SSCDW phase exhibits the same finite-size effect as the MCDW phase where one has to measure its properties with one boson less than the expected, then we use the same procedure from Sec. 3.2 for the calculation of the bosonic gaps, density profiles, and order parameters.. In Fig. 7(b), we show $\mathcal{O}^{\rm SSCDW}_{\sigma,\rho=2/5}$ for a range of $V_{\rm FF}$ values with the parameters of the balanced case from Fig. 7(a) for which we use either $\sigma=\uparrow$ or $\downarrow$ and $\rho_{\rm F}^{\sigma}=\rho_{\rm F}/2$. We note that after the critical value ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=2/5}\approx 0.1$ both the order parameter and the corresponding gap (Fig. 7(a)) increase from zero. For an imbalanced case, we checked that the definition (11) also works as an order parameter using the corresponding coupled fermionic component. Inside the balanced SSCDW phase, half of the fermions are uncoupled from the insulator structure, this decreases its stability with higher $V_{\rm FF}$ as seen in a reduction of both the gap and order parameter in Fig. 7. For large enough $V_{\rm FF}$, we expect this behavior to break the insulator and transition the ground state to a SF phase. Even then, the SSCDW phase appears only for $\rho_{\rm{B}}+\rho_{\rm{F}}\geq 1$ since in this region the chain gets saturated with all of the carriers, which forces the coupling to only contain one type of fermion and not both, in contrast to what happens with the MCDW. This condition is also independent of the spin-population imbalance since it only deals with the complete fermionic and bosonic densities.
@@ 3.4 Phase diagram for nonzero spin-population imbalance
According to the previous sections, there are two spin-selective phases in the present long-ranged mixture model, so it is of great value to analyze how the phase diagram from Fig. 1(a) changes as we increase the spin-population imbalance. This is depicted in Fig. 10, where we exhibit the phase diagram for $I=0$, $1/6$, $1/2$ and $1$ , using the corresponding density relations for non-zero spin-population imbalance found in previous sections.
@@
Here, we observe that the phase space configuration of the MMI (dark blue) and MCDW (brown) insulators is invariant under the change of the spin-population imbalance, which further emphasizes its spin-independent nature. On the other hand, as we increase $I$, we can see the splitting of the spin-selective phases. In the case of the SSMI (dark green), the separation grows until the spin-up insulator follows $\rho_{\rm{B}}=1$ while the other one matches the MMI line. For the SSCDW (light green), there is an additional constraint, $\rho_{\rm{B}}+\rho_{\rm{F}}\geq 1$; hence as the imbalance increases, the spin-down phase shortens until it vanishes at $I=1$, while the spin-up phase follows the relation $\rho_{\rm{B}}-\rho_{\rm{F}}=0$. In the case of the ICDW (purple line), the condition $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$ limits its phase space line as $I$ grows until it converges at the point $\rho_{\rm{B}}=\rho_{\rm{F}}=1/2$. Hence, for polarized fermions at half-filling the three different insulators found in this study follow the single condition $\rho_{\rm{B}}-\rho_{\rm{F}}=0$. This emphasizes how a study with spinor fermions can show different emergent behaviors that can be elusive in an investigation with only scalar particles.
@@ 4 Bosonic next-neighbor interactions
Up to this point, we have only considered nonzero next-neighbor fermionic interactions, then a good question is what happens when the long-ranged interaction is between the bosons only. In order to answer this, we plot in Fig. 11 the $\rho_{\rm{B}}-\mu_{\rm{B}}$ graph for $\rho_{\rm{F}}=2/5$, $2/3$, $4/5$, keeping the local interactions constant at $U_{\rm{FF}}=4$ and $U_{\rm{BF}}=6$ while turning on the next-neighbor bosonic interaction $V_{\rm{BB}}$ for values in the range $0\leq V_{\rm{BB}}\leq 6$.
@@
Figure 11: Bosonic density $\rho_{\rm{B}}$ vs. bosonic chemical potential $\mu_{\rm{B}}$ at the thermodynamic limit for fermionic densities (a) $\rho_{\rm{F}}=2/3$, (b) $\rho_{\rm{F}}=2/5$ and (c) $\rho_{\rm{F}}=4/5$. The local repulsion couplings are $U_{\rm{FF}}=4$, and $U_{\rm{BF}}=6$, and we compare the behavior with and without next-neighbor interaction between bosons using $V_{\rm{BB}}=0$, $2$, $3$, $6$. Each inset shows the charge gap $\Delta_{\rm{B}}$ as a function of the inverse of the system size $L$ for the corresponding nonzero $V_{\rm{BB}}$ at characteristic bosonic densities $\rho_{\rm{B}}$. The points correspond to DMRG results and the lines are visual guides.
@@
For $\rho_{\rm{F}}=2/3$ (Fig. 11(a)) and $V_{\rm{BB}}=2$ we find once again three plateaus corresponding to the MMI at $\rho_{\rm{B}}=1/3$, the SSMI at $\rho_{\rm{B}}=2/3$ and the ICDW at $\rho_{\rm{B}}=1/2$ with charge gaps $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/3}=1.27$, $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/3}=2.51$ and $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=1.18$, respectively. In this case, the main difference concerning $V_{\rm{FF}}\neq 0$ is that the plateaus of the MMI increase slightly, instead of decreasing as it happens in Fig. 4(a), which shows that the charge gap has a dependence on the associated particle statistics of the long-ranged interactions. Apart from that, for $\frac{1}{2}\leq\rho_{\rm{F}}\leq\frac{2}{3}$ we found that just as Fig. 11(a), the phase diagram for fermionic next-neighbor interactions is the same for bosonic next-neighbor interactions (See Fig. 1).
@@
On the other hand, at $\rho_{\rm{F}}=2/5$ (Fig. 11(b)) we do observe an additional incompressible phase besides the ones in Fig. 4(a) when we turn on the long-ranged interaction to $V_{\rm{BB}}=6$. This new plateau is located at $\rho_{\rm{B}}=1/2$ and corresponds to the ICDW with charge gap $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=6.64$. Here, since the bosons possess long-ranged interactions they can form the CDW structure even in the absence of fermions, then in this case the ICDW appears for any fermionic filling, as can be seen in Fig. 1(b). Apart from this insulator, we still find the MMI and SSMI at $\rho_{\rm{B}}=3/5$ and $\rho_{\rm{B}}=4/5$ with gaps $\Delta_{\rm{B}}^{\rho_{\rm{B}}=3/5}=2.09$ and $\Delta_{\rm{B}}^{\rho_{\rm{B}}=4/5}=2.31$, respectively. For this instance, both of them increase their stability with higher $V_{\rm{BB}}$. In Sec. 3.2 for $V_{\rm{FF}}\neq 0$ we found that the long-ranged interactions reduced drastically the spin-selective phase gap since the free fermions are enhanced by this interaction, and hence can disturb the insulating structure. In this case, only the bosons have long-ranged interactions, and because of that the stability of the phase is not damped. The MCDW phase still appears in the phase diagram but with a tiny charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/5}=0.26$, considering the high value of the bosonic interaction in this case, this suggests that the MCDW structure is more favorable for $V_{\rm{FF}}\neq 0$.
@@
Finally, in Fig. 11(c) we show the case for $\rho_{\rm{F}}=4/5$, which has analogous results as the previous two fermionic densities. The ICDW appears with a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=3.16$, as expected from Fig. 1(b). For the MMI and SSMI, their charge gaps increase to $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/5}=1.16$ and $\Delta_{\rm{B}}^{\rho_{\rm{B}}=3/5}=2.88$, respectively, the same way as it happened for Fig. 11(a). On the contrary, the SSCDW arises with a small charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/5}=0.26$ similar to the MCDW in Fig. 11(b).